# Generative-Contrastive Heterogeneous Graph Neural Network

### v1

@@ 4.2. Meta-path View
In this part, model input are multi-views constructed from different meta-paths. For example, in Figure 3 (a), we connect two nodes of the same type using the specified meta-path ‘PAP’. Thus, such a graph composed of nodes of the same type is referred to as a view based on the meta-path ‘PAP’. Based on preliminaries, we define a set of meta-paths as $\left\{\rho_{1},\ldots,\rho_{j},\ldots,\rho_{M}\right\}$, and the corresponding views are $\left\{G^{\rho_{1}},G^{\rho_{2}},\ldots,G^{\rho_{M}}\right\}$. $M$ is the number of meta-paths we have selected. Here, the model defines a graph $\mathcal{G}=(\mathcal{V},\boldsymbol{A},\boldsymbol{X})$. For each graph, $\mathcal{V}$ is the node set, $N=$ $|\mathcal{V}|$ is the number of nodes, $\mathbf{A}\in\{0,1\}^{N\times N}$ is the adjacency matrix, and $\mathbf{X}\in\mathbb{R}^{N\times d}$ is the input node feature matrix.
@@
where $p$ is the mask ratio we set. Uniform random sampling helps the encoder avoid potential bias centers, which refers to the bias arising when all neighbors of a node are either completely masked or completely visible. We mask the entire node features in the set $\mathcal{V}^{[Mask]}$ to obtain $\widetilde{\boldsymbol{X}}$.
@@
where $\widetilde{\boldsymbol{X}}$ is feature input with the $[Mask]$, and $\boldsymbol{H}\in\mathbb{R}^{N\times d}$ denotes code after the encoder. Next, to further encourage the decoder to learn compressed representations, we use a new ratio and the same sampling strategy to mask a different set of nodes $\mathcal{V}^{[ReMask]}$. We use a GNN with the same structure as the encoder to act as the decoder. However, the decoder can be GAT (Velickovic et al., 2017), GCN (He et al., 2020a), and its layer number and type can be independent of the encoder’s.
@@
where $\gamma_{\rho_{j}}$ denotes the weight of meta-path $\rho_{j}$. According to the attention mechanism, when the semantic attention score $\gamma_{\rho_{j}}$ is higher, it indicates the meta-path $\rho_{j}$ gets more importance in downstream tasks. We aggregate feature embeddings from all meta-paths to obtain the final node features under the meta-path view.
@@ 4.3. Hierarchical Contrastive learning
The model employs two sampling strategies. For intra-contrast, it uses a node-level contrast, whereas for inter-contrast, it adopts an edge-level contrast. Edge-level sampling is illustrated in Figure 4. Next, we get an edge sampling set $\mathcal{E}$.
@@
where $\lambda$ is a balance coefficient to control the weight of each view pair. We employed intra-contrast and inter-contrast for hierarchical contrastive learning and conducted analyses on their operational mechanisms. The intra-contrast assumes that meta-paths are interconnected and aim to maximize their mutual information. This method integrates the different semantics across meta-paths to form a generalized knowledge representation for downstream tasks. On the other hand, the inter-contrast approach attempts to enhance node representation by contrasting local and global information, preventing nodes from being influenced by noise from higher-order neighbors. The hierarchical contrast and sampling strategies are designed to construct a useful discriminator to enhance the adversarial interplay between the generative and contrastive models.
@@ 4.4. Model Optimization
Node classification. First, for the node classification task (Wang et al., 2019, 2021, 2023), the common practice is to train embeddings for all nodes using self-supervised loss. Then, we fine-tune the model on different datasets using a few labeled samples and a single-layer MLP. The node-level and edge-level strategies we propose typically lead to considerable loss differences in the initial phase. Therefore, it is necessary to balance their weights during the training phase. The loss generated by hierarchical contrastive learning is as follows:
@@
Time complexity analysis. The network schema view’s time complexity is $O\left(|V|\cdot D_{n}d^{2}\right)$, where $|V|$ denotes the edge number of a heterogeneous graph. $D_{n}$ is the average node degree and $d$ is the feature dimension. For the meta-path view, GC-HGNN has $M$ meta-path-based subgraphs. So, the time complexity of the meta-path view is $O\left(|\mathcal{V}|\cdot 2MLDd\right)$, where $L$ is the number of GNN layers. Then, the self-supervised loss contains $\lambda_{Intra}$ and $\lambda_{Inter}$. The time complexity to get positive samples and calculate $\lambda_{Intra}$ is $O\left(|\widetilde{\mathcal{V}}|\cdot d\right)$. For the $\lambda_{Inter}$, the complexity is $O\left(|\mathcal{V}|^{2}\cdot d\right)$. Because the edge set $\mathcal{E}$ is a matrix $\mathcal{E}\in\{0,1\}^{N\times N}$ composed of target nodes. In the link prediction task, the complexity of the additional $\mathcal{L}_{{bpr}}$ is $O\left(|\mathcal{O}|\cdot d\right)$. Thus, the overall time complexity of GC-HGNN is $O\left(|V|\cdot D_{n}d^{2}+|\mathcal{V}|\cdot 2MLD_{m}d+|\widetilde{\mathcal{V}}|\cdot d+|\mathcal{V}|^{2}\cdot d+|\mathcal{O}|\cdot d\right)$. This complexity is the time complexity for each epoch during training, and the overall training time will be determined by the number of training epochs.
@@ 5. EXPERIMENT
Last.fm 31,469 913,216 4 Link Prediction
@@ 5.1.1. Datasets
The performance of GC-HGNN has been validated on six real-world datasets. These datasets encompasses four datasets for node classification: ACM, DBLP, Aminer, and Freebase, as well as four datasets for link prediction: Last.fm, Yelp, Douban Movie and Douban Book. A statistical overview of all the datasets is presented in Table 5.
@@
  • ACM (Han et al., 2022). The target nodes are papers, which are divided into three classes. For each paper, there are 3.33 authors averagely, and one subject.
  • DBLP (Fu et al., 2020). The target nodes are authors, which are divided into four classes. For each author, there are 4.84 papers averagely.
  • Freebase (Tian et al., 2023). The target nodes are movies, which are divided into three classes. For each movie, there are 18.7 actors, 1.07 directors and 1.83 writers averagely.
  • Aminer (Wang et al., 2021). The target nodes are papers. We extract a subset of original dataset, where papers are divided into four classes. For each paper, there are 2.74 authors and 8.96 references averagely.
@@
  • Last.fm [^1] https://www.last.fm/. The music recommendation dataset contains three types of nodes: user, artist, and tag, along with four edge relationships. We predict whether there is an interaction between a user and an artist.
  • Yelp [^2] https://www.yelp.com/. The business recommendation dataset contains five types of nodes, including user, business, and city, along with five edge relationships. We predict whether there is an interaction between a user and an business.
  • Douban Movie [^3] https://m.douban.com/. The movie recommendation dataset contains six types of nodes, including user, movie, and actor, as well as six edge relationships. We predict the interactions between a user and a movie.
  • Douban Book (Shi et al., 2018). The book recommendation dataset contains seven types of nodes, including user, book, and author, as well as seven edge relationships. We predict the interactions between a user and a book.
@@ 5.4. Ablation Study(RQ3)
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{mae}$: This variant replaces the masked autoencoder (MAE) of the meta-path view with a normal GAT (Velickovic et al., 2017) encoder.
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{hcl}$: This variant cancels the hierarchical contrastive learning (HCL) module and directly uses the InfoNCE loss for downstream tasks.
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{m\&h}$: The third variant simultaneously removes the masked autoencoder (MAE) and hierarchical contrastive learning (HCL) modules.
As shown in Figure 5, it can be observed that, in all cases, removing the HCL module results in worse performance compared to other variants. Additionally, variants where mae and HCL are removed generally exhibit the worst performance in most cases. This further highlights two important points: 1) The graph structural information is overemphasized, and reconstructing the graph links is not necessary; 2) The existing heterogeneous contrastive learning strategies are not sufficient to serve as good discriminators to enhance the masked autoencoder effectively.
@@ 5.5. Interpretability and Visualization (RQ4)
In this study, we focus on discussing the various modules of GC-HGNN, attempting to provide a rational concept or explanation. We investigated the advantages of generative-contrastive model, and then further illustrate the motivation through the representation visualization.
Interpretability. Generative-contrastive representation learning (Wu et al., 2023; Yang et al., 2023; Goodfellow et al., 2020) utilizes discriminative loss functions as objectives, aiming to address some inherent shortcomings of generative reconstruction, such as sensitive and conservative distribution and low-level abstraction objective. In contrastive learning (Chen et al., 2020; Wang et al., 2021; He et al., 2020b; Wu et al., 2021), we can alleviate these issues by setting up appropriate discriminators to learn ‘distinguishable’ information for different samples. However, the encoder and decoder brought by generative learning (Devlin et al., 2018; Hou et al., 2022; Tian et al., 2023; Chowdhery et al., 2023) endow it with powerful embedding expression capabilities. The decoder requires the representation to be reconstructable, meaning that it contains all the information of the input. Furthermore, generative-contrastive self-supervised methods abandon the point-to-point reconstruction objective and instead pursue the alignment of distributions in the data. In summary, adversarial methods (Liu et al., 2021) absorb the advantages of both generative and contrastive methods, but also face the challenge of how to build a framework in the field of heterogeneous graphs. We propose a masked autoencoder to enhance the expressive power of meta-path perspectives, and then use hierarchical contrastive learning and designed sampling strategies to serve as enhanced discriminators. Experiments demonstrate that this is an effective and novel exploration.
Visualization. To more intuitively perceive the embeddings of different models, we use t-SNE (Van der Maaten and Hinton, 2008) to reduce the dimensionality of the embeddings and visualize them with labels. As shown in Figure 6, we select three representative models: unsupervised MP2vec, contrastive HeCo, and generative HGMAE. We can observe that MP2vec performs the worst, with different types of nodes mixed together and difficult to distinguish. HeCo can effectively differentiate categories, but each block is separated and does not form obvious clusters. HGMAE uses a hidden dimension of 256 and is a generative method, thus exhibiting more evident clustering phenomena. However, our GC-HGNN is based on a contrastive framework and demonstrates both clustering and uniform distribution, balancing the advantages of contrastive and generative approaches. This phenomenon also indicates that our generative-contrastive paradigm is effective.
@@ 5.6. Parameter Sensitivity (RQ5)
In this study, we explored the impact of key hyperparameters on heterogeneous graph node classification. We investigated the effects of hierarchical contrast coefficient, embedding dimension and positive samples.
@@ 6. Conclusion
In this paper, we propose a novel generative-contrastive heterogeneous graph neural network named GC-HGNN. It is a generative learning enhanced contrastive paradigm and a meaningful exploration. We combine meta-paths and masked autoencoders to create augmented contrastive views without disrupting the data structure. Then, we propose position-aware and semantic-aware sampling strategies to alleviate the synthesize hard negatives problem in heterogeneous graphs. Finally, we propose a hierarchical contrastive learning self-supervised framework to capture local and global information. In particular, the hierarchical contrastive learning and sampling strategies constitute the enhanced discriminator under this paradigm. Experiments demonstrate that this discriminator effectively improves the stability of output embeddings. The generative-contrastive framework significantly improves the performance of node classification and link prediction tasks. We compare our model with seventeen state-of-the-art models on eight real-world datasets. Our model outperforms the latest contrastive and generative baselines on multiple tasks.

### v2

@@ 4.2. Meta-path View
In this part, model inputs are multi-views constructed from different meta-paths. To mitigate the imbalanced data augmentation, we enhance multiple views through a generative mask autoencoder. For example, in Figure 3 (a), we connect two nodes of the same type using the specified meta-path ‘PAP’. Thus, such a graph composed of nodes of the same type is referred to as a view based on the meta-path ‘PAP’. Based on preliminaries, we define a set of meta-paths as $\left\{\rho_{1},\ldots,\rho_{j},\ldots,\rho_{M}\right\}$, and the corresponding views are $\left\{G^{\rho_{1}},G^{\rho_{2}},\ldots,G^{\rho_{M}}\right\}$. $M$ is the number of meta-paths we have selected. Here, the model defines a graph $\mathcal{G}=(\mathcal{V},\boldsymbol{A},\boldsymbol{X})$. For each graph, $\mathcal{V}$ is the node set, $N=$ $|\mathcal{V}|$ is the number of nodes, $\boldsymbol{A}\in\{0,1\}^{N\times N}$ is the adjacency matrix, and $\mathbf{X}\in\mathbb{R}^{N\times d}$ is the input node feature matrix.
@@
where $p$ is the mask ratio we set. The uniform random sampling (Hou et al., 2022) helps the encoder avoid potential bias centers, which refers to the bias arising when all neighbors of a node are either completely masked or completely visible. We mask the entire node features in the set $\mathcal{V}^{[Mask]}$ to obtain $\widetilde{\boldsymbol{X}}$.
@@
where $\widetilde{\boldsymbol{X}}$ is feature input with the $[Mask]$, and $\boldsymbol{H}\in\mathbb{R}^{N\times d}$ denotes code after the encoder. Next, to further encourage the decoder to learn compressed representations, we use a new ratio and the same sampling strategy to mask a different set of nodes $\mathcal{V}^{[ReMask]}$. We use a GNN with the same structure as the encoder to act as the decoder. However, the decoder can be GAT (Veličković et al., 2018), GCN (He et al., 2020a), and its layer number and type can be independent of the encoder’s.
@@
where $\gamma_{\rho_{j}}$ denotes the weight of meta-path $\rho_{j}$. According to the attention mechanism (Wang et al., 2019), when the semantic attention score $\gamma_{\rho_{j}}$ is higher, it indicates the meta-path $\rho_{j}$ gets more importance in downstream tasks. We aggregate feature embeddings from all meta-paths to obtain the final node features under the meta-path view.
@@ 4.3. Hierarchical Contrastive learning
The inputs for this part are the embeddings obtained from network schema and meta-path: $\boldsymbol{h}_{e}^{1}$ and $\boldsymbol{h}_{e}^{2}$. To alleviate sampling bias and information loss in the contrastive discriminator, we propose novel sampling strategies and hierarchical contrast. The model employs two sampling strategies. For intra-contrast, it uses a node-level contrast, whereas for inter-contrast, it adopts an edge-level contrast. Edge-level sampling is illustrated in Figure 4. Next, we get an edge sampling set $\mathcal{E}$.
@@
where $\lambda$ is a balance coefficient to control the weight of each view pair. We employed intra-contrast and inter-contrast for hierarchical contrastive learning and analyzed their operational mechanisms. The intra-contrast assumes that meta-paths are interconnected and aim to maximize their mutual information (Liu et al., 2021). This method integrates the different semantics across meta-paths to form a generalized knowledge representation for downstream tasks. On the other hand, the inter-contrast approach attempts to enhance node representation by contrasting local and global information, preventing nodes from being influenced by noise from higher-order neighbors. The hierarchical contrast and sampling strategies are designed to construct a useful contrastive discriminator to enhance the adversarial interplay between the generative and contrastive models.
@@ 4.4. Model Optimization
Node classification. First, the common practice for the node classification task (Park et al., 2020; Wang et al., 2021) is to train embeddings for all nodes using self-supervised loss. Then, we fine-tune the model on different datasets using a few labeled samples and a single-layer MLP. The node-level and edge-level strategies we propose typically lead to considerable loss differences in the initial phase. Therefore, it is necessary to balance their weights during the training phase. The loss generated by hierarchical contrastive learning is as follows:
@@
Time complexity analysis. The network schema view’s time complexity is $O\left(|V|\cdot D_{n}d^{2}\right)$, where $|V|$ denotes the edge number of a heterogeneous graph. $D_{n}$ is the average node degree and $d$ is the feature dimension. For the meta-path view, GC-HGNN has $M$ meta-path-based subgraphs. So, the time complexity of the meta-path view is $O\left(|\mathcal{V}|\cdot 2MLDd\right)$, where $L$ is the number of GNN layers. Then, the self-supervised loss contains $\lambda_{Intra}$ and $\lambda_{Inter}$. The time complexity to get positive samples and calculate $\lambda_{Intra}$ is $O\left(|\widetilde{\mathcal{V}}|\cdot d\right)$. For the $\lambda_{Inter}$, the complexity is $O\left(|\mathcal{V}|^{2}\cdot d\right)$. Because the edge set $\mathcal{E}$ is a matrix $\mathcal{E}\in\{0,1\}^{N\times N}$ composed of target nodes. In the link prediction task, the complexity of the additional $\mathcal{L}_{{bpr}}$ is $O\left(|\mathcal{O}|\cdot d\right)$. Thus, the overall time complexity of GC-HGNN is $O\left(|V|\cdot D_{n}d^{2}+|\mathcal{V}|\cdot 2MLD_{m}d+|\widetilde{\mathcal{V}}|\cdot d+|\mathcal{V}|^{2}\cdot d+|\mathcal{O}|\cdot d\right)$. This complexity is the time complexity for each epoch during training, and the number of training epochs will determine the overall training time.
@@ 5. EXPERIMENT
Last.fm 31,469 913,216 3 Link Prediction
@@ 5.1.1. Datasets
The performance of GC-HGNN has been validated on six real-world datasets. These datasets encompass four datasets for node classification: ACM, DBLP, Aminer, and Freebase, as well as four datasets for link prediction: Last.fm, Yelp, Douban Movie and Douban Book. A statistical overview of all the datasets is presented in Table 5.
@@
  • ACM (Han et al., 2022). The target nodes are papers (3 classes), averagely linked to 3.33 authors and one subject.
  • DBLP (Fu et al., 2020). The target nodes are authors (4 classes), averagely linked to 4.84 papers.
  • Freebase (Tian et al., 2023). The target nodes are movies (3 classes), averagely linked to 18.7 actors, 1.07 directors, and 1.83 writers.
  • Aminer (Wang et al., 2021). The target nodes are papers (4 classes), averagely linked to 2.74 authors and 8.96 references.
@@
  • Last.fm [^1] https://www.last.fm/ predicts user-artist interactions in a music recommendation setting, with node types including user, artist, and tag.
  • Yelp [^2] https://www.yelp.com/ predicts user-business interactions in a business recommendation setting, with node types including category, city, etc.
  • Douban Movie [^3] https://m.douban.com/ predicts user-movie interactions in a movie recommendation setting, with node types including actor, etc.
  • Douban Book (Shi et al., 2018) predicts user-book interactions in a book recommendation setting, with node types including author, etc.
@@ 5.4. Ablation Study(RQ3)
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{mae}$: This variant replaces the meta-path view’s masked autoencoder (MAE) with a normal GAT (Veličković et al., 2018) encoder.
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{sam}$: This variant replaces the sampling strategy (SAM) of GC-HGNN with random sampling while maintaining the same sampling ratio.
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{hcl}$: The third cancels the hierarchical contrastive learning (HCL) module and directly uses the InfoNCE (2010infonce) loss for downstream tasks.
As shown in Figure 5, it can be observed that, in all cases, removing the HCL module results in worse performance compared to other variants. Additionally, variants where SAM and HCL are removed generally exhibit the worse performance in most cases. This further highlights two important points: 1) The graph structural information is overemphasized, and reconstructing the graph links is not necessary; 2) The existing heterogeneous contrastive learning strategies are not sufficient to serve as good contrastive discriminators to enhance the masked autoencoder effectively.
@@ 5.5. Interpretability and Visualization (RQ4)
In this study, we focus on discussing the various modules of GC-HGNN, attempting to provide a rational concept or explanation. We investigated the advantages of the generative-contrastive model and then further illustrated the motivation through the representation visualization.
Interpretability. Generative-contrastive representation learning (Wu et al., 2023; Yang et al., 2023; Goodfellow et al., 2020) utilizes discriminative loss functions as objectives, aiming to address some inherent shortcomings of generative reconstruction, such as sensitive and conservative distribution and low-level abstraction objectives. In contrastive learning (Chen et al., 2020; Wang et al., 2021; He et al., 2020b; Wu et al., 2021), we can alleviate these issues by setting up appropriate contrastive discriminators to learn ‘distinguishable’ information for different samples. However, the encoder and decoder brought by generative learning (Kenton and Toutanova, 2019; Hou et al., 2022; Tian et al., 2023; Chowdhery et al., 2023) endow it with powerful embedding expression capabilities. The decoder requires the representation to be reconstructable, meaning that it contains all the information of the input. Furthermore, generative-contrastive self-supervised methods abandon the point-to-point reconstruction objective and instead pursue the alignment of distributions in the data. In summary, adversarial methods (Liu et al., 2021) absorb the advantages of both generative and contrastive methods, but also face the challenge of how to build a framework in the field of heterogeneous graphs. We propose a masked autoencoder to enhance the expressive power of meta-path perspectives, and then use hierarchical contrastive learning and designed sampling strategies to serve as enhanced contrastive discriminators. Experiments demonstrate that this is an effective and novel exploration.
Visualization. To more intuitively perceive the embeddings of different models, we use t-SNE (Van der Maaten and Hinton, 2008) to reduce the dimensionality of the embeddings and visualize them with labels. As shown in Figure 6, we select three representative models: unsupervised MP2vec (Dong et al., 2017), contrastive HeCo (Wang et al., 2021), and generative HGMAE (Tian et al., 2023). We can observe that MP2vec performs the worst, with different types of nodes mixed together and difficult to distinguish. HeCo can effectively differentiate categories, but each block is separated and does not form obvious clusters. HGMAE uses a hidden dimension of 256 and is a generative method, thus exhibiting more evident clustering phenomena. However, our GC-HGNN is based on a contrastive framework and demonstrates both clustering and uniform distribution, balancing the advantages of contrastive and generative approaches. This phenomenon also indicates that our generative-contrastive paradigm is effective.
@@ 5.6. Parameter Sensitivity (RQ5)
In this study, we explored the impact of key hyperparameters on heterogeneous graph node classification. Moreover, we investigated the effects of hierarchical contrast coefficient, embedding dimension and mask ratio.
@@ 6. Conclusion
In this paper, we propose a novel generative-contrastive heterogeneous graph neural network named GC-HGNN. It is a generative learning enhanced contrastive paradigm and a meaningful exploration. We combine meta-paths and masked autoencoders to create augmented contrastive views without disrupting the data structure. Then, we propose position-aware and semantic-aware sampling strategies to alleviate the synthesize hard negatives problem in heterogeneous graphs. Finally, we propose a hierarchical contrastive learning self-supervised framework to capture local and global information. In particular, the hierarchical contrastive learning and sampling strategies constitute the enhanced contrastive discriminator under this paradigm. Experiments demonstrate that this contrastive discriminator effectively improves the stability of output embeddings. The generative-contrastive framework significantly improves the performance of node classification and link prediction tasks. We compare our model with seventeen state-of-the-art models on eight real-world datasets. Our model outperforms the latest contrastive and generative baselines on multiple tasks.